a satellite is to be put into orbit 500km above the earths surface.if its vertical velocity after launching is 2000m/s at this height calculate the magnitude and direction of the impulse required to put the satellite directly into its orbit,if its mass is 50kg
1
Expert's answer
2019-06-24T09:30:09-0400
"F=m \times a=\frac{m \times v^2}{R}"

"F=G \times \frac{m \times Mearth}{R^2}"

Equalize these equations and find v:


"v= (\frac{G \times M}{R})^\frac{1}{2}= 7550 meter/sec"

Required speed to enter orbit 500 kilometers, at an initial velocity of 2000 meters per second, it will not enter a given orbit